OpenStudy (anonymous):

the business analyst of a major company notices that the profit made by the company in 2008 is exactly twice that made in 2009. The profit made in 2010 is r5million more than the profit made in 2009. If the total profit made by the company from 2008 to 2010 is r73 million, determine the profit made in 2008, 2009, 2010

OpenStudy (anonymous):

can someone plz help

OpenStudy (anonymous):

i'm here to help

OpenStudy (anonymous):

you're from india? i noticed the "r" = rupee?

OpenStudy (anonymous):

and of course, your name :]

OpenStudy (anonymous):

from mauritius in the indian ocean :) a small paradise island

OpenStudy (anonymous):

that's beautiful!!!

OpenStudy (anonymous):

i need to show the working for this question..but it seems so confusing

OpenStudy (anonymous):

you must have rich parents?

OpenStudy (anonymous):

well mauritius is great living here..i really need ur help for this quest..its for my dissertation

OpenStudy (anonymous):

okay let x=2008 y=2009 z=2010

OpenStudy (anonymous):

@nishta, create an equation that defines the year 2009 as the base profit. for example, if year2009 was x, then year 2008 would be 2x. then total these up to equal 73mill, which is the total profit of the three years

OpenStudy (anonymous):

x=2y 2y+5M=z 73M=x+y+z

OpenStudy (anonymous):

now you just plug it in

OpenStudy (anonymous):

@Tai, that is a bit more than needed...

OpenStudy (anonymous):

73M=x+(x/2)+(x/2)+5M

OpenStudy (anonymous):

x+(2x/2)+5M=73M

OpenStudy (anonymous):

2x+5M=73M

OpenStudy (anonymous):

x=34M

OpenStudy (anonymous):

you find the y and z values ;)

OpenStudy (anonymous):

@nishta -- if year 2009 is x, what would year 2010 be?